OLYMPIAD TRAINING SESSION 2002 Trinity College, Cambridge, 4-7 April

INFORMATION FOR PARTICIPANTS

Here is some information about the Training Session. If you would like more information, or are unsure about anything, please do not hesitate to contact me by telephone or e-mail (details later).

Expenses

We will reimburse all reasonable travel expenses (which usually means Standard Class rail fare, with any appropriate discounts). Please bring receipts with you, and also try to get a receipt if you take a taxi. Accommodation is provided free, as are all meals.

How to get to Trinity

We will be staying at Burrell’s Field in Grange Road, highlighted on the map. The railway station (also highlighted) is quite a distance away – if you are walking, allow about 40 minutes. However, if you have luggage then it might be better to take a taxi. If you are being dropped off by car, please note that it is now impossible to park on Grange Road – use Adams Road or Clarkson Road instead.

Arrival at Burrell’s Field

Enter through the main gate and proceed to the Porters’ Lodge, which is on your right inside the gate. The Porters will tell you how to get to your room, and where we are meeting at 2pm. If possible, try to arrive between 1pm and 1.30pm: this will give you time to find your room, get unpacked and relax. We will start promptly at 2pm.

Departure from Burrell’s Field

We should finish work by 3.30pm, when there will be a farewell cup of tea. If you are being picked up by car, suggest a pick-up time of around 4pm (at Grange Road).

What to bring with you

(i) Pens, pencils, rubber, ruler, compasses, protractor, large pad of A4 paper, and an A4-size clipboard (if you have one). (ii) An A4-size file or folder (essential for safe storage of the large amount of paper you will collect). (iii) This document, the map, the list of participants, the programme, and the Mathematical Olympiad Handbook. (iv) A reliable alarm clock. Your room is likely to have hand towels, soap and a shaver point. However, you might bring spares, just in case. Please note that bath towels are not provided.

Dress

Dress is as informal as you please. There will be the odd occasion when a shirt, tie and jacket would be appropriate for men and a smartish dress for ladies. For instance, there are two events hosted by Trinity College (the Trinity Talk and Prizes on Friday and the Reception on Saturday) at which you should be (at least moderately) smartly dressed.

List of Staff

Most of the people listed as ‘Staff’ will be present throughout. We also hope to have some ‘Old Olympians’ (members of previous IMO teams) dropping in to help from time to time. RA Richard (Mr Richard Atkins, Oundle School, Deputy Leader 1999-2003) CB Christopher (Dr Christopher Bradley, Clifton College, Bristol, Former Deputy Leader) CCF Ceri (Miss Ceri Fiddes, research student, University of Bath) TG Toby (Mr Toby Gee, research student, University of London, former team member) IBL Imre (Dr Imre Leader of Trinity College Cambridge, Chief Co-ordinator (i.e. head of marking) IMO 2002, former Leader and team member) ACM Adam (Prof Adam McBride of the University of Strathclyde, Academic Director of IMO 2002, and Chairman of the IMO Jury 2002, former Leader) DM David (Dr David Monk, University of Edinburgh, former Leader and Deputy Leader). VdS Vin (Dr Vin de Silva, postdoctoral student, University of Stanford, former Deputy Leader and team member) ASa Adrian (Mr Adrian Sanders, research student, University of Cambridge, former team member) GCS Geoff (Dr Geoff Smith of the Universities of Bath and Paris VI, Leader 2002, bon viveur)

The programme

All events will take place in Burrell’s Field, with the exception of breakfast, lunch, dinner and the Trinity Talk and Prizes, which take place in the main part of the college.

The aim of the Ice-Breaker is to loosen everyone up and allow people to get to know each other, and to eradicate any remaining traces of humour as we turn ourselves into manic drones, driven by patriotism and the quest for precious metal.

For the rest of the sessions on Thursday, and all of the sessions on Friday, students will work on problems in groups of 4 – the makeup of these groups will be announced at the beginning. In addition, each group will have a ‘Background Problem’ to work on collectively during the three days and present on Sunday afternoon.

On Saturday there will be 3 topics in parallel during each period, with 6 or 7 students going to each (and working individually). Allocation of students to topics will take place on Friday.

The First Selection Test is a 4\frac{1}{2}-hour exam, consisting of 3 questions, which will help us to select the 8-person squad (from which a team of 6 and two reserves will eventually be chosen) for this year’s IMO.

Contact

If someone needs to contact you while you are at the Training Session, they should phone Trinity College (01223-33-8400) and ask to be put through to the Burrell’s Field Porters’ Lodge, where they can leave a message for you.

Before the event you can contact me (Geoff Smith) in France at +33 01 45 87 83 65 during most evenings (please do not telephone after 8pm UK time except in emergency). After March 29 but before the Trinity Camp I will be based in Bath (01225 835660). Email sent to G.C.Smith@bath.ac.uk follows me round wherever I am, so please always use that address and ignore any address whence I might send e-mail to you. You can also contact the Deputy Leader, Richard Atkins, by telephone ( home 01473-736-232) or e-mail (ra@oundle.northants.sch.uk).